OpenStudy (anonymous):

The area of a given hexagon is equal to the area of an equilateral triangle whose perimeter is 36 inches. Find the length of a side of the regular hexagon.

OpenStudy (anonymous):

3a=36 a=12 area of equilateral triangle=36sqrt3 area of hexagon=(3 sqrt(3) s^2)/2 (3 sqrt(3) s^2)/2=36sqrt3 s^2=72/3=24 s=2sqrt6 \[side of hexagon=2\sqrt6\]